Whited Products recently completed a 4-for-1 stock split. Prior to the split, its stock sold for $120 per share. If the firm's total market value increased by 5% as a result of increased liquidity caused by the split, what was the stock price following the split?


$28.43
$29.93
$31.50
$33.08
$34.73
Answer #1

stock price = (120+0.05*120)*(1/4)

=126/4 = $31.50